OpenStudy (amurciano2):

How would you integrate 13 sec^4x with b=pi/4 and a=-I/4

OpenStudy (jchick):

what are a and b here?

OpenStudy (jchick):

we need to find the integration for sec ^4 x

OpenStudy (jchick):

Let u=tan(x)

OpenStudy (jchick):

Then du=(tan(x))′dx

OpenStudy (jchick):

=sec^2(x)dx

OpenStudy (jchick):

we have that sec^2(x)dx=du

OpenStudy (jchick):

so we can write the expression as

OpenStudy (amurciano2):